# Date of publication xxxx 00, 0000, date of current version xxxx 00, 0000. xxxx

### v2

@@ V-A MNIST Classification
V-A1 Results for M=5M{=}5 epochs
Training both models for only M=5M{=}5 epochs across N=50N{=}50 runs reveals statistically significant differences in both training and test accuracy. This indicates that (ReLU, Sine, 6:2) indeed converges faster than ReLU-only (Figure 13, da10.py):
@@ V-A1 Results for M=5M{=}5 epochs
CNNA (ReLU, Sine): 0.9957 (Std Dev: 0.0004)
CNNB (ReLU): 0.9939 (Std Dev: 0.0003)
p-value: 0.0000
The difference is statistically significant.
@@
CNNA (ReLU, Sine): 0.9910 (Std Dev: 0.0010)
CNNB (ReLU): 0.9905 (Std Dev: 0.0014)
p-value: 0.0280
The difference is statistically significant.
V-A2 Results for M=30M{=}30 epochs
When training both models for M=30M{=}30 epochs (again repeated N=50N{=}50 times), the difference in final accuracy — both training and test — disappears (Figure 14, da11.py). As in earlier experiments classifying noisy sinusoid and pure noise, CNNA (ReLU, Sine, 6:2) converges more rapidly during the early epochs, but the two models eventually reach similar accuracy levels.
Examining the training losses, both models appear to reach similar final performance. However, when we examine the test losses, both models reach a minimum and then begin to worsen, most likely due to overfitting. However, overfitting appears to be less severe in the (ReLU, sine, 6:2) model than in the pure ReLU model.
We therefore conclude that, in both the synthetic 1D signal classification task and the 2D MNIST image classification problem, networks equipped with a dual activation layer (ReLU, sine) converge more rapidly than networks using ReLU alone.
Figure 14: Training and test accuracies and losses over M=30M{=}30 epochs for MNIST classification, averaged over N=50N{=}50 runs. Comparison between (ReLU, sine, 6:2) and ReLU-only.
V-B ECG-ID Classification
@@
The dataset was partitioned into two stratified folds. For each run, one fold was used for training and the remaining fold was split evenly into validation and test sets. In this work, we report performance based on the validation set, omitting a separate test set as the focus remains on comparative training dynamics.
V-B1 Without batch normalization
We implemented a PyTorch program that follows the same experimental procedure used for MNIST classification (Section V-A), using the following architecture:
@@ V-B1 Without batch normalization
Training both models for M=500M{=}500 epochs across N=10N{=}10 runs reveals statistically significant differences in both training and test accuracy for both short and long training regimes (M=50M{=}50 and M=500M{=}500). These results show that, in this more complex problem, (ReLU, Sine, 1:1) not only converges faster but also achieves higher final training and test accuracy compared to pure ReLU (Figure 15, ecgid.ipynb).
Figure 15: Evolution of training and validation accuracies and losses over M=500M{=}500 epochs for ECG-ID classification, averaged over N=10N{=}10 runs, without batch normalization. The model using (ReLU, Sine, 1:1) activations converges faster and attains statistically better performance than the pure ReLU baseline.
@@
CNNA (ReLU, Sine): 0.3147 (Std Dev: 0.0292)
CNNB (ReLU): 0.1555 (Std Dev: 0.0285)
p-value: 0.0000
@@
CNNA (ReLU, Sine): 0.3178 (Std Dev: 0.0246)
CNNB (ReLU): 0.1580 (Std Dev: 0.0239)
p-value: 0.0000
@@
CNNB (ReLU): 0.9678 (Std Dev: 0.0123)
p-value: 0.0000
@@
CNNA (ReLU, Sine): 0.9372 (Std Dev: 0.0114)
CNNB (ReLU): 0.8483 (Std Dev: 0.0211)
p-value: 0.0000
@@
V-B2 With batch normalization
Figure 16: Comparison between (ReLU, sine, 1:1) and ReLU-only models, both including a batch normalization after the second convolutional block. Training was conducted for M=500M{=}500 epochs and repeated over N=10N{=}10 independent runs.
Integrating a batch normalization layer after the second convolutional block and prior to average pooling significantly diminishes the performance gap between the (ReLU, sine, 1:1) and ReLU-only models on training data. Nevertheless, the DAL model (ReLU, sine, 1:1) consistently maintains its superiority in validation performance, both during the initial stages (M=50M{=}50) and upon reaching full convergence (M=500M{=}500).
@@ V-B2 With batch normalization
CNNA (ReLU, Sine): 0.7915 (Std Dev: 0.0373)
CNNB (ReLU): 0.7616 (Std Dev: 0.0575)
p-value: 0.0413
@@
CNNA (ReLU, Sine): 0.6655 (Std Dev: 0.0504)
CNNB (ReLU): 0.6108 (Std Dev: 0.0674)
p-value: 0.0084
The difference is statistically significant.
@@
The difference is not statistically significant.
@@
CNNA (ReLU, Sine): 0.9355 (Std Dev: 0.0145)
CNNB (ReLU): 0.9054 (Std Dev: 0.0208)
p-value: 0.0000
@@
V-B3 Dead neurons without and with batch normalization
Table II reports the percentage of dead neurons for models without batch normalization, whereas Table III shows the corresponding values with batch normalization. The results indicate that networks using (ReLU, Sine, 1:1) are notably less prone to neuron inactivation. Without normalization, the ReLU-only model exhibits severe neuron death — reaching up to 71% in dense layers — while the dual-activation model remains considerably more stable.
Batch normalization substantially reduces neuron death in both architectures. Once normalization is applied, the training curves of the two models become nearly identical (Figure 16), suggesting that batch normalization compensates for much of ReLU’s inherent instability. Even so, the (ReLU, sine, 1:1) model consistently maintains lower dead-neuron rates, supporting the idea that combining ReLU with sine activations improves gradient flow compared to using ReLU alone.
@@ V-B3 Dead neurons without and with batch normalization
Layer (ReLU, sine, 1:1) Only ReLU
@@
conv3 8.34% 33.51%
dense1 25.00% 70.84%
dense2 7.15% 39.29%
@@
Layer (ReLU, sine, 1:1) Only ReLU
@@ VI Conclusions
Our results further reveal that the benefit of sine does not arise from periodicity, but rather from its local non-zero derivative shape near x=0x{=}0. Activation functions such as tanh and truncated sine — despite being non-periodic — produce similar improvements when mixed with ReLU. We found that typical activation inputs in our networks lie within the interval [1.5,+1.5][-1.5,+1.5], meaning that the behavior of the activation near zero plays a dominant role during training.
A key finding of this work is that mixing ReLU with a smooth, zero-centered activation such as sine helps mitigate the dying ReLU problem, reducing the number of inactive neurons and enabling more stable gradient propagation. This improved gradient flow leads to faster convergence, even in architectures that do not suffer from strictly zero gradients, such as those using GELU, Swish, or Leaky ReLU.
Based on these insights, we proposed the DAL, which combines two activation functions in a single layer — for example, ReLU and sine. Across a variety of tasks, including synthetic 1D signal classification, MNIST digit recognition, and ECG-ID signal classification, DAL consistently accelerates convergence and often reaches final performance slightly better than that of standard activations. The computational overhead introduced by the sine function is minimal, ensuring DAL operates with nearly the same efficiency as a standard ReLU layer.

### v3

@@ VI-A MNIST Classification
VI-A1 MNIST classification for M=5M{=}5 epochs
Given the near-ceiling performance levels on this task, a ceiling effect obscures the performance gap, making it difficult to determine whether one algorithm statistically outperforms the other. Nevertheless, training both models for only M=5M{=}5 epochs over N=50N{=}50 runs reveals statistically significant disparities: the differences in training and test accuracy yielded $p$-values of p<0.001p{<}0.001 and p=0.0280p{=}0.0280, respectively. This indicates that DAL indeed converges faster than ReLU-only (Figure 13, da10.py).
@@ VI-A1 MNIST classification for M=5M{=}5 epochs
CNNA (ReLU, Sine): 0.9957
CNNB (ReLU): 0.9939
One-sided (>) paired p-value: 0.0000
@@
CNNA (ReLU, Sine): 0.9910
CNNB (ReLU): 0.9905
One-sided (>) paired p-value: 0.0280
VI-A2 MNIST classification for M=30M{=}30 epochs
Even when training both models for M=30M{=}30 epochs over N=50N{=}50 repetitions, the disparity in final training and test accuracy persists and remains statistically significant (p=0.0020p{=}0.0020 and p=0.0165p{=}0.0165).
Training Accuracy Comparison:
CNNA (ReLU, Sine): 0.9994
CNNB (ReLU): 0.9991
One-sided (>) paired p-value: 0.0020
Test Accuracy Comparison:
CNNA (ReLU, Sine): 0.9923
CNNB (ReLU): 0.9918
One-sided (>) paired p-value: 0.0165
When we examine the test losses (Figure 14), both models reach a minimum and then begin to worsen, most likely due to overfitting. However, overfitting appears to be less severe in the DAL model than in the pure ReLU model.
Figure 14: Training and test accuracies and losses over M=30M{=}30 epochs for MNIST classification, averaged over N=50N{=}50 runs. Comparison between (ReLU, sin, 6:2) and ReLU-only.
VI-B ECG-ID Classification
@@
The dataset was partitioned into two stratified folds. For each run, one fold was used for training and the remaining fold was used to evaluate.
VI-B1 ECG-ID Classification Without Batch Normalization
We implemented a PyTorch program that follows the same experimental procedure used for MNIST classification (Section VI-A), using the following architecture:
@@ VI-B1 ECG-ID Classification Without Batch Normalization
Training both models for M=500M{=}500 epochs across N=10N{=}10 runs reveals statistically significant differences in both training and test accuracy for both short and long training regimes (M=50M{=}50 and M=500M{=}500). Note that given the 90-class nature of the dataset, the baseline for random chance is approximately 1.1%, placing the observed accuracy in its proper context. These results show that, in this more complex problem, the DAL model converges faster and achieves higher final training and test accuracy compared to pure ReLU (Figure 15, ecgid.ipynb).
Figure 15: Evolution of training and test metrics for ECG-ID classification. Results are averaged over N=10N{=}10 independent runs for M=500M{=}500 epochs without batch normalization. The DAL architecture (ReLU, sin, 1:1) exhibits accelerated convergence and achieves statistically significant performance gains compared to the pure ReLU baseline.
@@
CNNA (ReLU, Sine): 0.3146 (Std Dev: 0.0264)
CNNB (ReLU): 0.1459 (Std Dev: 0.0244)
One-sided (>) paired p-value: 0.0000
@@
CNNA (ReLU, Sine): 0.3107 (Std Dev: 0.0268)
CNNB (ReLU): 0.1535 (Std Dev: 0.0246)
One-sided (>) paired p-value: 0.0000
@@
CNNB (ReLU): 0.9629 (Std Dev: 0.0125)
One-sided (>) paired p-value: 0.0000
@@
CNNA (ReLU, Sine): 0.9379 (Std Dev: 0.0093)
CNNB (ReLU): 0.8507 (Std Dev: 0.0206)
One-sided (>) paired p-value: 0.0000
@@
The experiments in this section were implemented in PyTorch using the library’s default weight initialization. To ensure robustness, we also evaluated the models by explicitly enforcing both Glorot (Xavier) uniform and He (Kaiming) uniform initializations across all layers. In all test scenarios — for both M=50M{=}50 and M=500M{=}500 epochs — the DAL model achieved statistically significant higher accuracy than the ReLU-only baseline.
VI-B2 ECG-ID Classification With Batch Normalization
Figure 16: Comparison between (ReLU, sin, 1:1) and ReLU-only models, both including a batch normalization after the second convolutional block. Training was conducted for M=500M{=}500 epochs and repeated over N=10N{=}10 independent runs.
Integrating a batch normalization layer after the second convolutional block and prior to average pooling accelerates the convergence of both networks (compare Figures 15 and 16). Under this configuration, the training accuracy curves for both models become nearly identical, suggesting that batch normalization compensates for much of ReLU’s inherent instability. Regarding test performance, DAL consistently outperforms ReLU-only; while this lead is not statistically significant in the early stages, it achieves statistical significance once full convergence is attained.
@@ VI-B2 ECG-ID Classification With Batch Normalization
CNNA (ReLU, Sine): 0.8058 (Std Dev: 0.0397)
CNNB (ReLU): 0.7721 (Std Dev: 0.0569)
One-sided (>) paired p-value: 0.0125
@@
CNNA (ReLU, Sine): 0.6581 (Std Dev: 0.0506)
CNNB (ReLU): 0.6229 (Std Dev: 0.0896)
One-sided (>) paired p-value: 0.0768
The difference is not statistically significant.
@@
p-value cannot be computed.
@@
CNNA (ReLU, Sine): 0.9321 (Std Dev: 0.0145)
CNNB (ReLU): 0.8968 (Std Dev: 0.0208)
One-sided (>) paired p-value: 0.0000
@@
VI-B3 Dead neurons without and with batch normalization
Table II reports the percentage of dead neurons for models without batch normalization, whereas Table III shows the corresponding values with batch normalization. The results indicate that networks using DAL(ReLU, sin, 1:1) are notably less prone to neuron inactivation. Without normalization, the ReLU-only model exhibits severe neuron death — reaching up to 75% in dense layers — while the dual-activation model remains considerably more stable.
Batch normalization substantially reduces neuron death in both architectures. Even so, the DAL model consistently presents lower dead-neuron rates, higher test accuracy and lower test loss, supporting the idea that combining ReLU with sine activations improves gradient flow compared to using ReLU alone.
@@ VI-B3 Dead neurons without and with batch normalization
Layer (ReLU, sin, 1:1) Only ReLU
@@
conv3 8.16% 33.86%
dense1 23.34% 75.00%
dense2 7.15% 41.67%
@@
Layer (ReLU, sin, 1:1) Only ReLU
@@ VII Conclusions
Our results further reveal that the benefit of sine does not arise from periodicity, but rather from its local non-zero derivative shape near x=0x{=}0. Activation functions such as tanh and truncated sine — despite being non-periodic — produce similar improvements when mixed with ReLU.
A key finding of this work is that mixing ReLU with a smooth, zero-centered activation such as sine helps mitigate the dying ReLU problem, reducing the number of inactive neurons and enabling more stable gradient propagation. This improved gradient flow leads to faster convergence, even in architectures that do not suffer from strictly zero gradients, such as those using GELU, Swish, or Leaky ReLU. Based on these insights, we proposed the DAL, which combines two activation functions in a single layer — for example, ReLU and sine.
When comparing two architectures of equal size, the DAL model demonstrates a higher density of active units than its conventional counterparts. Consequently, adopting the DAL approach enhances the network’s effective width, providing the benefits of a larger architecture without the computational cost of additional parameters.
Across a diverse range of tasks — including synthetic 1D signal classification, MNIST digit recognition, and ECG-ID biometric recognition — the DAL consistently accelerates convergence. It frequently yields a final performance superior to standard activations, with these improvements reaching statistical significance in the majority of cases. Furthermore, the computational overhead introduced by the sine function is marginal, ensuring that the DAL maintains nearly the same efficiency as a conventional ReLU layer.